# User Modeling and User Profiling: A Comprehensive Survey

### 15 Feb 2024 v1

@@ 3.7 Deep learning
The advent of deep neural network-based models has sparked a revolutionary shift across various fields, including user modeling. These models have significantly contributed to advancements in this field, enabling more accurate and comprehensive user profiling and prediction of user behavior. Different deep learning approaches and architectures have been employed in user modeling. In particular, differentiable user models (e.g., [Hämäläinen et al., 2023]), attention mechanism (e.g., [Wang et al., 2020c, 2022a; Fazil et al., 2021; Qi et al., 2021, 2022b; Chu et al., 2022], graph neural networks (e.g., [Chen et al., 2019b, 2021b; Dai and Wang, 2021; Wu et al., 2021a; Yan et al., 2021, 2022; Han et al., 2022; Zhao et al., 2022; Cheng et al., 2023]), convolutional neural networks (e.g., [Wang et al., 2019b; Wanda and Jie, 2020; Fazil et al., 2021; Qi et al., 2022b]), autoencoders (e.g., [Abu Sulayman and Ouda, 2019; Wang et al., 2019b; Fazelnia et al., 2022; Liu et al., 2023c]), recurrent neural networks (e.g., [Ge et al., 2018; Ni et al., 2018; Yu et al., 2019; Gu et al., 2020; Chu et al., 2022; Li et al., 2022b]), long-short term memory (e.g., [Żołna and Romański, 2017; Ma et al., 2020; Fazil et al., 2021; Sahoo and Gupta, 2021; Nkambou et al., 2023]), and transformers (e.g., [Gu et al., 2021a; Kota et al., 2021; Zhu et al., 2021; Avny Brosh et al., 2022; Wu et al., 2022; Zheng et al., 2022a]).
@@ 4.2.5 User profile representation
Semantic network-based (Ontology-based)
@@ 4.2.6 Modeling techniques
  • Ontology-based modeling concerns using ontologies, which define concepts and relationships in a domain, to represent and reason about user preferences, characteristics, and behavior. It utilizes ontological rules for making inferences about users. Rimitha et al. [2018] presented a method for enhancing personalized job recommender systems through the development of ontology-based user profiles. It underscores the significance of user profiles in delivering tailored job suggestions to individuals by accurately reflecting user preferences and experiences. Chari et al. [2020] introduced the Explanation Ontology, a comprehensive model specifically crafted to encapsulate the various facets of explanations within user-centered AI systems. The model is presented as a valuable tool for system designers, offering a structured approach to user modeling with ontology in the context of AI explanations. It emphasizes the importance of user-centric design in AI and the ongoing efforts to update and refine the ontology to support the dynamic landscape of user needs and system capabilities.
@@
  • Collaborative filtering relies on user-item interactions and recommendations from a group of users with similar tastes. It can be user-based, item-based, or a combination of both. Ma et al. [2022a] investigated the effects of rapid, population-scale concept drifts, such as those induced by pandemic-like events, on collaborative filtering models within Recommender Systems (RSs). The core of the study is the introduction of the NEST framework, which simulates the evolution of user needs and behaviors in response to non-stationary external events. Wang et al. [2022a] introduced a recommender system framework known as Time-aware Attention-based Deep Collaborative Filtering (TADCF), whose core is its ability to model dynamic user preferences over time.
@@
    • Boosting algorithms are a set of ensemble learning techniques that combine the predictions of multiple weak models to create a robust and accurate predictive model. These algorithms sequentially train weak learners, with each subsequent model focusing on the mistakes of its predecessors. Popular boosting algorithms include AdaBoost and Gradient Boosting, which are applied to capture complex relationships in user data. An extensive study on the application of boosting algorithms for user profiling, specifically for predicting student performance in higher education, has been presented by Hamim et al. [2022]. To achieve this, the analysis compares different research studies that utilize boosting algorithms for student profile modeling. The paper also explores the impact of student behaviors and characteristics on their academic performance. It proposes a student profile model that leverages educational data mining techniques to predict, classify, and adaptively support students’ learning processes. This model can also be used for e-recommender systems that tailor educational content to individual student needs.
@@
    • Dimensionality reduction techniques reduce the number of features in the dataset while retaining its essential information. Reducing dimensionality aids in visualizing user data patterns and improving computational efficiency. Principal Component Analysis (PCA) is the most popular method of this type. Bi et al. [2016] introduced an anomaly detection model to scrutinize user behavior within database systems and web browsing environments. PCA is utilized to analyze user behavior by calculating covariance matrices and extracting significant behavior features. This statistical approach simplifies the complexity of user behavior data, allowing the model to distinguish between normal and abnormal patterns effectively.
@@ 4.4.1 Recommender Systems
4.4.2 Personalized web search
Personalized search is a process that prioritizes a user’s interests when searching through a set of documents or web pages. This is achieved by creating a user profile based on the sites the user has searched for, thereby identifying their interests. The accuracy of the system is enhanced by predicting the user’s interests more accurately than a typical search. In these personalized search systems, the search results are ranked not only based on the query information but also according to the user’s interests. This approach ensures a more tailored and relevant search experience for the user. To cite specific cases, Zhou et al. [2020b] proposed a personalized search model that aims to improve the re-finding behavior of users when they search for previously encountered information. The model leverages memory networks to capture two distinct types of re-finding behavior: query-based and document-based. It utilizes Recurrent Neural Networks (RNNs) to process sequential data and generate refined vectors that represent user intent over multiple sessions. Zhou et al. [2021a] introduced a group-based model that employs both search behavior and friend networks to enhance search results for users, particularly those with sparse historical data. By grouping users into friend circles and utilizing neural networks, the model creates a more accurate semantic representation of user interests, leading to significant improvements in search personalization. Deng et al. [2022a] leveraged a dual-feedback network to enhance the understanding of user search intentions. The model is designed to improve personalized search by incorporating multi-granular user feedback, both positive and negative, to accurately model the user’s current search intention. Other works can be seen in Du et al. [2016], Ma et al. [2020], and Zhou et al. [2021b].

### 20 Feb 2024 v2

@@ 3.7 Deep learning
The advent of deep neural network-based models has sparked a revolutionary shift across various fields, including user modeling. These models have significantly contributed to advancements in this field, enabling more accurate and comprehensive user profiling and prediction of user behavior. Different deep learning approaches and architectures have been employed in user modeling. In particular, differentiable user models (e.g., [Hämäläinen et al., 2023]), attention mechanism (e.g., [Wang et al., 2020c, 2022a; Fazil et al., 2021; Qi et al., 2021, 2022b; Chu et al., 2022]), graph neural networks (e.g., [Chen et al., 2019b, 2021b; Dai and Wang, 2021; Wu et al., 2021a; Yan et al., 2021, 2022; Han et al., 2022; Zhao et al., 2022; Cheng et al., 2023]), convolutional neural networks (e.g., [Wang et al., 2019b; Wanda and Jie, 2020; Fazil et al., 2021; Qi et al., 2022b]), autoencoders (e.g., [Abu Sulayman and Ouda, 2019; Wang et al., 2019b; Fazelnia et al., 2022; Liu et al., 2023c]), recurrent neural networks (e.g., [Ge et al., 2018; Ni et al., 2018; Yu et al., 2019; Gu et al., 2020; Chu et al., 2022; Li et al., 2022b]), long-short term memory (e.g., [Żołna and Romański, 2017; Ma et al., 2020; Fazil et al., 2021; Sahoo and Gupta, 2021; Nkambou et al., 2023]), and transformers (e.g., [Gu et al., 2021a; Kota et al., 2021; Zhu et al., 2021; Avny Brosh et al., 2022; Wu et al., 2022; Zheng et al., 2022a]).
@@ 4.2.5 User profile representation
Semantic-based (Ontology-based)
@@ 4.2.6 Modeling techniques
  • Ontology-based modeling concerns using ontologies, which define concepts and relationships in a domain, to represent and reason about user preferences, characteristics, and behavior. It utilizes ontological rules for making inferences about users. Rimitha et al. [2018] presented a method for enhancing personalized job RSs through the development of ontology-based user profiles. It underscores the significance of user profiles in delivering tailored job suggestions to individuals by accurately reflecting user preferences and experiences. Chari et al. [2020] introduced the Explanation Ontology, a comprehensive model specifically crafted to encapsulate the various facets of explanations within user-centered AI systems. The model is presented as a valuable tool for system designers, offering a structured approach to user modeling with ontology in the context of AI explanations. It emphasizes the importance of user-centric design in AI and the ongoing efforts to update and refine the ontology to support the dynamic landscape of user needs and system capabilities.
@@
  • Collaborative filtering relies on user-item interactions and recommendations from a group of users with similar tastes. It can be user-based, item-based, or a combination of both. Ma et al. [2022a] investigated the effects of rapid, population-scale concept drifts, such as those induced by pandemic-like events, on collaborative filtering models within RSs. The core of the study is the introduction of the NEST framework, which simulates the evolution of user needs and behaviors in response to non-stationary external events. Wang et al. [2022a] introduced a recommender system framework known as Time-aware Attention-based Deep Collaborative Filtering (TADCF), whose core is its ability to model dynamic user preferences over time.
@@
    • Boosting algorithms are a set of ensemble learning techniques that combine the predictions of multiple weak models to create a robust and accurate predictive model. These algorithms sequentially train weak learners, with each subsequent model focusing on the mistakes of its predecessors. Popular boosting algorithms include AdaBoost and Gradient Boosting, which are applied to capture complex relationships in user data. An extensive study on the application of boosting algorithms for user profiling, specifically for predicting student performance in higher education, has been presented by Hamim et al. [2022]. To achieve this, the analysis compares different research studies that utilize boosting algorithms for student profile modeling. The paper also explores the impact of student behaviors and characteristics on their academic performance. It proposes a student profile model that leverages educational data mining techniques to predict, classify, and adaptively support students’ learning processes. This model can also be used for e-RSs that tailor educational content to individual student needs.
@@
    • Dimensionality reduction techniques reduce the number of features in the dataset while retaining its essential information. Reducing dimensionality aids in visualizing user data patterns and improving computational efficiency. Principal Component Analysis (PCA) is the most popular method of this type. Bi et al. [2016] introduced an anomaly detection model to scrutinize user behavior within database systems and web browsing environments. PCA analyzes user behavior by calculating covariance matrices and extracting significant behavior features. This statistical approach simplifies the complexity of user behavior data, allowing the model to distinguish between normal and abnormal patterns effectively.
@@ 4.4.1 Recommender systems
4.4.2 Information retrieval
In the information retrieval context, user modeling research focuses particularly on personalized search, which is a process that prioritizes a user’s interests when searching through a set of documents or web pages. This is achieved by creating a user profile based on the sites the user has searched for, thereby identifying their interests. The accuracy of the system is enhanced by predicting the user’s interests more accurately than a typical search. In these personalized search systems, the search results are ranked not only based on the query information but also according to the user’s interests. This approach ensures a more tailored and relevant search experience for the user. To cite specific cases, Zhou et al. [2020b] proposed a personalized search model that aims to improve the re-finding behavior of users when they search for previously encountered information. The model leverages memory networks to capture two distinct types of re-finding behavior: query-based and document-based. It utilizes Recurrent Neural Networks (RNNs) to process sequential data and generate refined vectors that represent user intent over multiple sessions. Zhou et al. [2021a] introduced a group-based model that employs both search behavior and friend networks to enhance search results for users, particularly those with sparse historical data. By grouping users into friend circles and utilizing neural networks, the model creates a more accurate semantic representation of user interests, leading to significant improvements in search personalization. Deng et al. [2022a] leveraged a dual-feedback network to enhance the understanding of user search intentions. The model is designed to improve personalized search by incorporating multi-granular user feedback, both positive and negative, to accurately model the user’s current search intention. Other works can be seen in Du et al. [2016], Ma et al. [2020], and Zhou et al. [2021b].